The shape of the distribution of the time required to get an oil change at a 20​-minute ​oil-change facility is unknown.​ However, records indicate that the mean time is 21.6 minutes​, and the standard deviation is 4.1 minutes. Complete parts c ​(c) Suppose the manager agrees to pay each employee a​ $50 bonus if they meet a certain goal. On a typical​ Saturday, the​ oil-change facility will perform 35 oil changes between 10 A.M. and 12 P.M. Treating this as a random​ sample, there would be a​ 10% chance of the mean​ oil-change time being at or below what​ value? This will be the goal established by the manager. There is a​ 10% chance of being at or below a mean​ oil-change time of nothing minutes. ​(Round to one decimal place as​ needed.)
Answer #1
mean μ= 21.6
standard deviation σ= 4.1000
sample size       =n= 35
std error=σ=σ/√n= 0.693
for 10th percentile critical value of z= -1.28
therefore corresponding value=mean+z*std deviation= 20.7 minute